john invested 50,000, part of it at 6% and the other part at 8%. The annual interest on the 6% investment was 480 more than that from the 8% investment. How much was invested at each rate?
Answer by Boreal(15235)

x=amount at 6%
50000-x=amount at 8%
0.06x-480=0.08(50000-x)=4000-0.08x
move the -0,08x to the left and the 480 to the right
0.14x=4480
x=$32,000@6%=$1920
50,000-x=$18,000@8%=$1440, and that is $480 less than $1920.